Enter your answer in the box provided. What is the molality of a solution consisting of 31.9 mL of benzene (C6H6; d = 0.877 g/mL) in 343 mL of hexane (C6H14; d = 0.660 g/mL)
Answer #1

moality = mol solute / kg solvent

MW benzene= 78.11 G7mol

mol solute= mass/MW = D*V/MW = (31.9*0.877)/(78.11 ) = 0.358165 mol of S

kg solvnet = DV = 343*0.66/1000 = 0.22638 kg

then

molality = 0.358165/0.22638 = 1.58214molal